# CoT-Space: A Theoretical Framework for Internal Slow-Thinking via Reinforcement Learning

### v1

@@ Abstract
Reinforcement Learning (RL) has become a pivotal approach for enhancing the reasoning capabilities of Large Language Models (LLMs). However, a significant theoretical gap persists, as traditional token-level RL frameworks fail to align with the reasoning-level nature of complex, multi-step thought processes like Chain-of-Thought (CoT). To address this challenge, we introduce CoT-Space, a novel theoretical framework that recasts LLM reasoning from a discrete token-prediction task to an optimization process within a continuous, reasoning-level semantic space. By analyzing this process from both a noise perspective and a risk perspective, we demonstrate that the convergence to an optimal CoT length is a natural consequence of the fundamental trade-off between underfitting and overfitting. Furthermore, extensive experiments provide strong empirical validation for our theoretical findings. Our framework not only provides a coherent explanation for empirical phenomena such as overthinking but also offers a solid theoretical foundation to guide the future development of more effective and generalizable reasoning agents.
@@ 1 Introduction
Reasoning is a critical capability for Large Language Models (LLMs), and there has been a significant focus on developing methods to improve it. One of the most well-known frameworks is Chain-of-Thought (CoT) [1]. Under the CoT framework, a promising approach to enhance LLM reasoning is “test-time scaling”, also known as “slow thinking” [2, 3]. This branch of methods generally falls into two categories: external and internal. External methods extend the LLM’s output by introducing additional steps, often through a tree-structured search [4]. These techniques, which do not require extra training [5], include popular frameworks such as Monte-Carlo Tree Search (MCTS)[6, 7] and self-consistency (also known as Best-of-N) [8]. In contrast, internal methods integrate the slow-thinking capability directly into the model itself through post-training. This is achieved either by supervised fine-tuning (SFT), a process often referred to as distillation, or by reinforcement learning (RL).
@@
Recent advancements in reasoning models such as DeepSeek’s R1 [9], OpenAI’s o1 and o3 [10, 11], and Qwen’s QwQ [12] have intensified the focus on the internal slow thinking methods, and especially using RL to enhance LLM reasoning. This training paradigm is often termed zero, indicating that the model is post-trained directly from its pre-training checkpoint using RL, bypassing the SFT stage. The results of zero-like approach are encouraging, as they demonstrate the potential for LLMs to develop reasoning capabilities through self-exploration with minimal supervision. This observation aligns with the core philosophy of RL: by setting a high-level, global objective, the policy can be guided to discover an optimal solution independently. As shown in Figure 1, the application of RL to LLM reasoning shares similarities with classical RL scenarios. For instance, as shown in subgraph (a), in a classic RL task like Bricks Breaker [^1] Bricks Breaker is a video game, the goal is to destroy the bricks by shooting a ball at them., the agent’s goal is to break more bricks. Without explicit instructions on specific strategies, the agent may discover a highly effective method, such as hitting the ball behind the top row of bricks to achieve a higher score. This kind of emergent behavior is a common outcome in classical RL. Similarly, in the context of LLM reasoning, as shown in subgraph (b), the policy model is given the broad goal of providing a correct answer. As RL training progresses, the model converges on an effective CoT pattern, demonstrating its ability to learn complex reasoning pathways from a simple, high-level reward signal.
Building on the discovery that models can automatically converge on a suitable CoT, various RL-based methods have been proposed for more effective zero-like post-training, including GRPO [13], VC-PPO [14], DAPO [15], and VAPO [16]. However, despite these practical advances, a significant gap remains in the theoretical understanding of LLM internal “slow thinking” via RL. Firstly, the surprising “overthinking” phenomenon [17] suggests there is an optimal CoT length for the reasoning process. Although empirical observations show different convergence patterns with various RL recipes, the precise mechanism driving this behavior remains unknown. Secondly, the generalization ability of LLM reasoning tasks is not yet well understood; while we observe this phenomenon, the underlying principles are not fully analyzed.
A central challenge to the theoretical analysis of LLM reasoning via RL is the misalignment between the traditional RL theoretical framework and the CoT paradigm. The conventional framework for RL with LLMs operates at the token level, where each action is the selection and concatenation of a single token to the current output. In contrast, the CoT approach functions as a higher-level reasoning framework, where the output is generated step-by-step, with each thought containing multiple tokens to solve a specific reasoning step. To address this limitation, we propose CoT-Space, a novel reasoning-level theoretical framework for LLM reasoning via RL. This framework is designed to enable a more precise theoretical analysis of this domain. We further validate its effectiveness by performing two corresponding analyses aimed at understanding both the convergence of CoT length and the generalization of reasoning tasks, thereby addressing the key theoretical gaps previously identified.
The remainder of this paper is structured as follows. In Section 2, we provide a systematic analysis of the misalignment between the current token-level RL formulation and the reasoning-level CoT framework. Subsequently, we introduce our reasoning-level theoretical framework, CoT-Space. We then leverage this framework in Section 3 to analyze the convergence of CoT length in the reasoning process from the perspective of noise and risk, demonstrating its potential theoretical value. We perform empirical validation of our theoretical insights in Section 4. Following this, we briefly review related works in Section 5 and finally conclude the paper in Section 6.
@@
Figure 1: An illustration of the analogy between strategy discovery in classical RL and RL for LLM reasoning. (a) In a classical RL scenario such as a Brick Breaker game, the agent is given a high-level goal to break more bricks. Through exploration, it discovers a non-trivial, effective strategy to hit the ball behind the bricks to maximize its reward. (b) In LLM reasoning, the policy model is given a high-level goal to provide the correct answers. As RL training progresses, the model autonomously discovers that generating a suitable CoT is an effective strategy to achieve the goal.
@@ 2 CoT-Space: A Reasoning-Level Theoretical Framework
This section presents a systematic analysis of the limitations inherent in the current token-level RL formulation when applied to the higher-order cognitive processes of CoT reasoning. We first discuss the fundamental misalignment of this conventional approach in Subsection 2.1 and then discuss the continuum nature of the reasoning-level state space in Subsection 2.2. Finally, we introduce our novel, reasoning-level theoretical framework CoT-Space in Subsection 2.3.
@@ 2.1 From Token-Level to Reasoning-Level Perspective
The conventional theoretical framework for applying RL to LLMs is fundamentally misaligned with the reasoning-level nature of CoT scenarios. This token-level assumption is inherited from traditional RL frameworks, where the agent receives a reliable and meaningful reward at each state transition. For instance, classic RL tasks like chess, Go, or video games allow for a meaningful measurement of a state’s value, which guides the agent’s policy.
However, in the context of natural language, the auto-regressive nature of modern language models has led us to formulate each token prediction as a state transition, resulting in a popular token-level theoretical framework. We briefly review this formulation here. Prior works usually model natural language generation as a token-level Markov Decision Process (MDP): $\mathcal{M}(\mathcal{S},\mathcal{A},r,H)$, where $S$ is the state space of existing token sequences, $A$ is the action space representing the token dictionary, $r$ is the reward function guiding generation, and $H$ is the token budget. A policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ produces the next token, and the state transitions by concatenating the predicted token $a=\pi(s)$ to the existing sequence $s$. This formulation presents a significant challenge regarding the state space $\mathcal{S}$ as the following assertion.
@@
On the one hand, the state space $\mathcal{S}$ can be partitioned into a union of infinite subspaces based on sequence length $t$: $\mathcal{S}=\mathcal{S}_{0}\cup\mathcal{S}_{1}\cup\cdots\cup\mathcal{S}_{t}\cup\cdots$, where $\mathcal{S}_{t}=\mathcal{A}^{t}$. As the combinations of dictionary tokens grow exponentially, the subspace for a given prompt $q$, $\mathcal{S}_{|q|}$, and all subsequent subspaces $\mathcal{S}_{>|q|}$ are astronomically large. Consequently, the data observed during RL training constitutes only a minute fraction of the entire huge possible token-level state space. On the other hand, the transition between states is unidirectional and discrete due to the auto-regressive nature of the process: $\forall\bm{s}_{t}\in\mathcal{S}_{t},\mathcal{T}(\bm{s}_{t},\pi)\in\mathcal{S}_{t+1}$. This indicates the discrete nature of the token-level state space.
Turning to CoT reasoning processes, we observe that a complete reasoning trajectory of length $L$ comprises multiple conceptual steps $(\xi_{1},\xi_{2},\dots,\xi_{L})$ with each step $\xi_{l}$ realized through multiple token generations. Consider solving algebraic equations: each $\xi_{l}$ corresponds to a specific subgoal in the solution process, achieved through coordinated token generation. This perspective reveals two fundamentally distinct procedural layers:
  1. 1.
    Reasoning-Level: The strategic planning of correct reasoning steps to solve problems
  2. 2.
    Token-Level: The tactical execution of individual reasoning steps through token generation
Traditional MDP formulations reduce LLM reasoning to token-level exploration, while the essence of reasoning involves higher-order cognitive processes that orchestrate sequences of conceptual steps. The token-level perspective fails to capture the strategic planning and structural coherence inherent in human-like reasoning, creating substantial theoretical barriers to understanding modern reasoning paradigms such as CoT.
Thus, to establish theoretically sound foundations, we must transition from token-level formulations to reasoning-level perspectives that properly reflect the problem’s intrinsic nature.
@@ 2.2 On the Continuum Nature of the Reasoning-Level State Space
Building upon Assertion 2.1, which characterizes the token-level state space as both high-dimensional and discrete, we now provide a theoretical basis for approximating this space as a continuum at the reasoning level. The core insight is that while the token-level space consists of discrete points, the semantic space it represents becomes increasingly dense as the reasoning process unfolds.
The essence of reasoning lies in the semantic meaning of conceptual steps, not in the specific tokens used to express them. A single reasoning step, such as “calculating the total weekly pages” of the instance in Fig. 1, can be realized by a vast number of distinct token sequences:
  • “Step 2: He writes 6*2=12 pages every week.”
  • “Then, I’ll calculate the total weekly pages: 6 pages/friend * 2 friends = 12 pages.”
  • “So, for both friends, the page count per week is 6 $\times$ 2, which equals 12.”
These sequences correspond to distant, discrete states in the token-level space. However, in a higher-dimensional semantic space, they are nearly identical, occupying a very small neighborhood. The CoT-Space framework posits that as the CoT length $L$ increases, the number of semantically meaningful states grows so rapidly that the space they inhabit can be treated as a continuous manifold. This density arises from the expressive redundancy of language: the number of valid token-level paths to achieve a reasoning goal grows exponentially with the available token budget.
@@
(Semantic equivalence set.) For an abstract reasoning step $\xi_{l}$ (e.g., a calculation, a logical deduction), its semantic equivalence set $\mathcal{T}(\xi_{l})$ is the collection of all token sequences $\tau\in\mathcal{A}^{*}$ that are semantically equivalent to $\xi_{l}$.
$\mathcal{T}(\xi_{l})=\{\tau\in\mathcal{A}^{*}\ |\ \text{Decode}(\tau)\equiv\xi_{l}\},$
where $\mathcal{A}^{*}$ is the space of all possible token sequences and $\text{Decode}(\cdot)$ is an abstract function mapping a token sequence to its semantic meaning.
@@
$N_{\text{realize}}(\xi_{1},...,\xi_{L})=\prod_{l=1}^{L}|\mathcal{T}(\xi_{l})|.$
Definition 2.4.
(Reasoning state density.) Within a fixed semantic volume $\mathcal{V}_{\text{semantic}}$, the set of all valid, reachable reasoning states of reasoning token amount $K$ is denoted by $\mathcal{S}_{\text{reasoning}}^{(K)}$. The state density $\rho(L)$ is defined as the number of such states per unit of semantic volume.
$\rho(K)=\frac{|\mathcal{S}_{\text{reasoning}}^{(K)}|}{|\mathcal{V}_{\text{semantic}}|}.$
To substantiate our claim that this density grows exponentially, we propose the following intuitive assumption.
@@
(Exponential expressive redundancy of language.) For any non-trivial reasoning step $\xi_{l}$, the size of its semantic equivalence set $|\mathcal{T}(\xi_{l},k)|$, when realized using an average of $k$ tokens, grows exponentially with $k$. That is, there exists a constant $c>1$ such that:
$|\mathcal{T}(\xi_{l},k)|\geq c^{k}.$
This assumption is grounded in the combinatorial nature of language; allowing even a few extra tokens for modifiers, rephrasing, or synonyms creates an exponentially larger space of semantically equivalent expressions.
@@
(Continuum convergence of reasoning-level state space.) Under Assumption 2.5, the reasoning-level state density, denoted as $\rho(K)$, increases exponentially with the total reasoning token amount $K$. This exponential growth implies that the expected semantic distance between any two adjacent valid reasoning states decreases dramatically as $K$ increases, and approaches zero as $K\to\infty$. Therefore, in the macroscopic limit, the discrete reasoning-level state space converges to a continuous semantic manifold.
The proof is provided in Appendix A.1. Theorem 2.6 provides the formal justification for the central premise of our work. This result licenses our departure from the discrete, token-level MDP formulation and enables the adoption of more powerful, analytically tractable tools from continuous optimization theory. With the continuum nature of the reasoning space established, we are now positioned to formally introduce our framework. In the subsequent section, we will leverage this continuum perspective to define the CoT-Space and reframe the LLM reasoning process as an optimization problem within this newly characterized landscape.
2.3 Formulation of CoT-Space
Building upon the discussion above, which characterizes the token-level state space and illustrates the continuum nature of reasoning-level state space, we now develop a reasoning-level theoretical framework designed to bridge the analytical gap. This alternative perspective draws an analogy to interstellar travel in space, as shown in Figure 2. From a token-level perspective, each individual token and its corresponding state are like discrete points, or stars, in a vast semantic galaxy. The auto-regressive generation process then resembles a journey across these stars. Since transitions only occur between these discrete states, the generation process is fundamentally discrete at the token-level scale, as illustrated in panel (a).
However, from the reasoning-level perspective, a continuum approximation emerges when we “zoom out” to consider the limiting behavior of this discrete system. Due to the exponential density of semantically valid states, the global properties of the space converge toward continuum-like characteristics, as visualized in panel (b). This macroscopic viewpoint allows us to move beyond individual tokens and analyze reasoning from a more abstract, strategic level.
Figure 2: An illustration for the shift from a token-level to a reasoning-level perspective. (a) The traditional token-level scale views the generation process as a sequence of discrete transitions in a vast state space. (b) The proposed reasoning-level scale zooms out to view the state space from a macroscopic perspective, where the incredibly dense and numerous token-level states can be approximated as a continuous semantic reasoning space.
To enable this higher-level analysis, we provide the following definitions for a new formalization of LLM reasoning via RL. We begin by defining the reasoning-level state. Based on the CoT setting, where the output is generated step-by-step and consists of multiple intermediate thoughts, a reasoning state is defined as follows.

### v2

@@ Abstract
Reinforcement Learning (RL) has become a pivotal approach for enhancing the reasoning capabilities of Large Language Models (LLMs). However, a significant theoretical gap persists, as traditional token-level RL frameworks fail to align with the reasoning-level nature of complex, multi-step thought processes like Chain-of-Thought (CoT). To address this challenge, we introduce CoT-Space, a novel theoretical framework that recasts LLM reasoning from a discrete token-prediction task to an optimization process within a continuous, reasoning-level semantic space. This shift in perspective serves as a conceptual bridge, revitalizing foundational principles from classical learning theory to analyze the unique dynamics of LLMs. By analyzing this process from both a noise perspective and a risk perspective, we demonstrate that the convergence to an optimal CoT length is a natural consequence of the fundamental trade-off between underfitting and overfitting. Furthermore, extensive experiments provide strong empirical validation for our theoretical findings. Our framework not only provides a coherent explanation for empirical phenomena such as overthinking but also offers a solid theoretical foundation to guide the future development of more effective and generalizable reasoning agents. We open-source our code at https://github.com/ZyGan1999/CoT-Space.
@@ 1 Introduction
Improving the reasoning capabilities of Large Language Models (LLMs) is pivotal in modern AI research. The most well-known framework for this is Chain-of-Thought (CoT) (Wei et al., 2022b), which elicits multi-step reasoning. A promising strategy to enhance CoT is “slow-thinking” (Chen et al., 2025b; Jiang et al., 2024), which can be broadly categorized into external and internal methods. External methods, such as tree-based strategies (Yao et al., 2023; Zhang et al., 2024; Wan et al., 2024) and self-consistency (Wang et al., 2022), augment the model’s inference-time output via searching (Gan et al., 2025). In contrast, internal methods embed slow-thinking directly into the model via post-training, typically using supervised fine-tuning (SFT) or reinforcement learning (RL).
Recent breakthroughs from models like DeepSeek’s R1 (DeepSeek-AI, 2025), OpenAI’s o1 and o3 (OpenAI, 2024; 2025), and Qwen’s QwQ (Qwen, 2024) have spotlighted the power of using RL for internal slow thinking. This paradigm, often called zero-like training, involves applying RL directly to a pre-trained model, bypassing SFT. The results are compelling, showing that LLMs can learn sophisticated reasoning strategies through self-exploration. This aligns with a core tenet of RL: a high-level objective can guide a policy to discover an optimal strategy on its own.
@@
As illustrated in Figure 1, RL for LLM reasoning is analogous to classical RL scenarios. In a classic RL task like Bricks Breaker [^1] Bricks Breaker is a video game, the goal is to destroy the bricks by shooting a ball at them. shown in subgraph (a), the agent’s goal is to break more bricks. Without explicit instructions on specific strategies, the agent may discover a highly effective method, such as hitting the ball behind the top row of bricks to achieve a higher score. This kind of emergent behavior is a common outcome in classical RL. Similarly, in the context of LLM reasoning, as shown in subgraph (b), the policy model is given the broad goal of providing a correct answer. As RL training progresses, the model converges on an effective CoT pattern, demonstrating its ability to learn complex reasoning pathways from a simple, high-level reward signal.
Building on this, a variety of RL-based methods have been developed to improve zero-like post-training, including GRPO (Shao et al., 2024), VC-PPO (Yuan et al., 2025), DAPO (Yu et al., 2025), and VAPO (Yue et al., 2025b). However, despite these empirical successes, our theoretical understanding of internal slow thinking remains shallow. For instance, the “overthinking” phenomenon (Sui et al., 2025), which suggests an optimal CoT length exists, lacks a clear mechanistic explanation. Furthermore, the principles governing the generalization of LLM reasoning are not yet well understood.
@@
Figure 1: Analogy between strategy discovery in classical RL and LLM reasoning. (a) In classical RL, an agent with a high-level goal (e.g., break more bricks) discovers an effective strategy through exploration to maximize its reward. (b) Similarly, an LLM policy with the goal of providing correct answers autonomously learns that generating a suitable CoT is an effective strategy.
This chasm between empirical success and theoretical understanding is alarming. It stands in stark contrast to classical machine learning, where decades of research have forged a robust theoretical bedrock. Foundational principles like optimization theory and generalization bounds, the very cornerstones of traditional ML, appear to falter when applied to the complex, multi-step nature of LLM reasoning. This prompts the critical question that motivates our work: Are these time-tested theories truly obsolete in this new era, or do we merely lack the conceptual framework to bridge them to the unique dynamics of LLM reasoning? We argue for the latter, positing that by shifting our analytical perspective, we can revitalize these foundational theories and use them to build a principled understanding of LLM reasoning.
The central barrier to this goal is twofold. First, a fundamental misalignment exists between token-level RL frameworks and the reasoning-level nature of CoT, where actions are complete thoughts rather than single tokens. Second, a conceptual gap separates the discrete world of language from the continuous mathematics underpinning classical learning theory, hindering the application of its powerful analytical tools. To resolve both challenges, we introduce CoT-Space, a novel reasoning-level theoretical framework for LLM reasoning via RL. Our framework first defines a reasoning-level state space to align RL with CoT’s structure. Subsequently, we prove this space converges to a continuous manifold, a key result that recasts reasoning as an optimization process. By leveraging this continuous perspective, we are then able to conduct analyses that explain the convergence of CoT length and the generalization of reasoning, revitalizing classical theory to address the key theoretical gaps previously identified.
The remainder of this paper is structured as follows. In Section 2, we provide a systematic analysis of the misalignment between the current token-level RL formulation and the reasoning-level CoT paradigm. Subsequently, we introduce our reasoning-level theoretical framework, CoT-Space. We then leverage this framework in Section 3 to analyze the convergence of CoT length in the reasoning process from the perspective of noise and risk, demonstrating its potential theoretical value. We perform empirical validation of our theoretical insights in Section 4. Following this, we briefly review related works in Section 5 and finally conclude the paper in Section 6.
@@ 2 CoT-Space: A Reasoning-Level Theoretical Framework
This section presents a systematic analysis of CoT reasoning via RL. We first discuss this misalignment (Subsection 2.1), then introduce our reasoning-level framework, CoT-Space (Subsection 2.2), and finally prove its continuum nature to establish its theoretical foundation (Subsection 2.3).
@@ 2.1 From Token-Level to Reasoning-Level Perspective
The conventional theoretical framework for RL in LLMs is fundamentally misaligned with the reasoning-level nature of CoT scenarios. This misalignment stems from modeling auto-regressive language generation as a token-level Markov Decision Process (MDP) (Setlur et al., 2025): $\mathcal{M}(\mathcal{S},\mathcal{A},r,H)$, where $\mathcal{S}$ is the state space of existing token sequences, $\mathcal{A}$ is the action space representing the token dictionary, $r$ is the reward function guiding generation, and $H$ is the token budget. In this standard formulation, a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ produces the next token $a=\pi(s)$ to extend the current sequence $s$. While inherited from classic RL, this approach presents a significant challenge regarding the token-level state space $\mathcal{S}$ in natural language, as we assert next.
@@
Consider a CoT reasoning process comprises multiple conceptual steps $(\xi_{1},\xi_{2},\dots,\xi_{L})$, which naturally reveals two distinct procedural layers: (1) Reasoning-Level: The strategic planning of abstract reasoning steps $(\xi_{1}\rightarrow\xi_{L})$. (2) Token-Level: The tactical realization of specific steps via token generation. Traditional MDP formulations, however, operate only at the problematic token level, failing to capture the higher-order planning inherent to CoT.
The token-level state space $\mathcal{S}$ is problematic for two reasons. First, it grows exponentially with sequence length ($\mathcal{S}_{t}=\mathcal{A}^{t}$). Second, the auto-regressive process imposes a discrete and unidirectional topology, further complicating analysis. This narrow focus creates a theoretical barrier that obscures the problem’s true nature. To establish a sound foundation, we must therefore transition from a token-level formulation to a reasoning-level perspective.
2.2 Formulation of CoT-Space
Figure 2: Token-level vs. Reasoning-level perspectives. (a) The token-level view treats the generation process as a path through discrete states. (b) The reasoning-level view zooms out, approximating the dense state space as continuous.
Building upon the discussion above, we now develop a reasoning-level theoretical framework designed to bridge the analytical gap. This alternative perspective draws an analogy to interstellar travel in space, as shown in Figure 2. From a token-level perspective, each individual token and its corresponding state are like discrete stars in a vast semantic galaxy. The auto-regressive generation process then resembles a journey across these stars. Since transitions only occur between states, the generation process is fundamentally discrete at the token-level scale, as illustrated in panel (a).
In contrast, from the reasoning-level perspective, a continuum approximation emerges when we “zoom out” to consider the limiting behavior of this discrete system. Due to the exponential density of semantically valid states, the global properties of the space converge toward continuum-like characteristics, as visualized in panel (b). This macroscopic viewpoint allows us to move beyond individual tokens and analyze reasoning from a more abstract, stepwise level.
To enable this higher-level analysis, we provide the following definitions for a new formalization of LLM reasoning via RL. Based on the CoT setting, where the output is generated step-by-step and consists of multiple intermediate thoughts, a reasoning state is defined as follows.
Definition 2.2.
(CoT reasoning state.) Given a query $q$ and an incomplete set of intermediate reasoning steps $\bm{\xi}_{o}=[\xi_{1},\xi_{2},\dots,\xi_{t}]$, the current reasoning state is defined as $s_{o}=(q,\bm{\xi}_{o})$.
Definition 2.2 establishes the concept of reasoning-level state. We then introduce a special class of states that represent successful problem solutions, which we call minimums.
Definition 2.3.
(Minimums in reasoning space.) Given a query $q$ with a definite golden answer $\phi$. Assume $\Xi_{q}$ is the set of all reasonable intermediate reasoning processes for query $q$. Then the minimums of the reasoning space are defined as $\bm{M}_{q}=\left\{(q,\bm{\xi},\phi)~|~\bm{\xi}\in\Xi_{q}\right\}$.
According to Definition 2.3, minimums are special states where the query has been successfully solved. For any intermediate reasoning state, we can then define which minimums are reachable.
Definition 2.4.
(Reachable minimums.) Given a query $q$ with a definite golden answer $\phi$, and an incomplete set of intermediate reasoning steps $\bm{\xi}_{o}=[\xi_{1},\xi_{2},\dots,\xi_{t}]$. The reachable minimums are defined as $\bm{R}^{o}_{q}=\left\{(q,\bm{\xi},\phi)~|~\bm{\xi}_{o}\sqsubseteq\bm{\xi}~\text{and}~\bm{\xi}\in\Xi_{q}\right\}$, where $a\sqsubseteq b$ represents that $a$ is prefix of $b$.
Definition 2.5.
(Reachable distance and nearest reachable minimum.) Given a query $q$ with a definite golden answer $\phi$, and a reasoning state $s_{o}$ on this query, the reachable distance between the $s_{o}$ and a reachable minimum is defined as $\operatorname{dist}(s_{o},m)=|\bm{\xi}|-|\bm{\xi}_{o}|$, where $m\in\bm{R}^{o}_{q}$, $\bm{\xi}$ and $\bm{\xi}_{o}$ are intermediate steps of $m$ and $s_{o}$ respectively, $|\cdot|$ represents the length of a vector. The nearest minimum is defined as $m^{*}_{o}={\arg\min}_{m\in\bm{R}^{o}_{q}}\,\operatorname{dist}(s_{o},m)$.
Definition 2.6.
(Reasoning loss.) Given a query $q$ with a definite golden answer $\phi$, and a reasoning state $s_{o}$ on this query, there exists a loss function $C$ that quantifies the distance of $s_{o}$ towards its nearest reachable minimum, which satisfies: for two different states $s_{i},s_{j}$, $C(s_{i})<C(s_{j})$ if and only if $\operatorname{dist}(s_{i},m^{*}_{i})<\operatorname{dist}(s_{j},m^{*}_{j})$. Specially, $\forall m\in\bm{M}_{q},C(m)=0$.
Definition 2.6 conceptualizes the reasoning process as an optimization process. The function $C(\cdot)$ acts as a proxy for the distance to a correct solution, where a lower loss indicates a more advanced stage of reasoning. Taking Figure 3(a) as instance, in this transition, $m_{i},m_{j}$ and $m_{k}$ are all reachable minimums for the current state $s_{0}$. However, the possible next states of $s_{0}$, which are denoted as $s_{i},s_{j}$ and $s_{k}$ (Here we refer to the three realizations $s_{i_{1}},s_{i_{2}}$ and $s_{i_{3}}$ all as $s_{i}$ for simplicity). These intermediate states differ in their distance to its corresponding minimums ($\operatorname{dist}(s_{i},m_{i})<\operatorname{dist}(s_{j},m_{j})<\operatorname{dist}(s_{k},m_{k})$), thus the reasoning losses satisfy $C(s_{i})<C(s_{j})<C(s_{k})$.
With this complete CoT-Space formulation, we resolve the misalignment between token-level RL and the reasoning-level CoT paradigm. We now turn to prove the continuum nature of reasoning-level state space, thus yielding the rationality of CoT-Space.
Figure 3: Illustration of CoT-Space. (a) Different reasoning steps ($\xi_{i},\xi_{j},\xi_{k}$) lead to different minimums ($m_{i},m_{j},m_{k}$), while one step can have multiple token-level realizations ($\xi_{i_{1}},\xi_{i_{2}},\xi_{i_{3}}$). (b) The discrete state space $\mathcal{S}$ (dots) is approximated as a continuous manifold $\tilde{\mathcal{S}}$ (blue area), where the real optimization vector ($\vec{v}_{\text{real}}$) approximates the ideal one ($\vec{v}_{\text{ideal}}$).
@@ 2.3 On the Continuum Nature of the Reasoning-Level State Space
The essence of reasoning lies in the semantic meaning of conceptual steps, not in the specific tokens used to express them. A single reasoning step, such as “calculating the total weekly pages” can be realized by a vast number of distinct token sequences $\xi_{i_{1}},\xi_{i_{2}}$ and $\xi_{i_{3}}$, as shown in Figure 3(a).
From token-level perspective, different token sequences correspond to distant, discrete states. However, in a higher-dimensional semantic space, they are nearly identical. The CoT-Space framework posits that as the reasoning length increases, the number of semantically meaningful states grows so rapidly that the space they inhabit can be treated as a continuous manifold. This density arises from the expressive redundancy of language: the number of valid token-level paths to achieve a reasoning goal grows exponentially with the available token budget.
@@
(Semantic equivalence set.) For an abstract reasoning step $\xi_{l}$ (e.g., a calculation, a logical deduction), its semantic equivalence set $\mathcal{V}(\xi_{l},k)$ is the collection of all token sequences $\tau\in\mathcal{A}^{k}$ with length $k$ that are semantically equivalent to $\xi_{l}$, defined as: $\mathcal{V}(\xi_{l},k)=\{\tau\in\mathcal{A}^{k}\ |\ \text{Decode}(\tau)\equiv\xi_{l}\},$ where $\mathcal{A}^{k}$ is the space of all possible token sequences with length $k$, and $\text{Decode}(\cdot)$ is an abstract function mapping a token sequence to its semantic meaning.
@@
(Reasoning state density.) Within a fixed semantic volume $\mathcal{V}_{\text{semantic}}$, the set of all valid, reachable reasoning states of reasoning token amount $K$ is denoted by $\mathcal{S}_{\text{reasoning}}^{(K)}$. The state density $\rho(K)$ is defined as the number of states per unit of semantic volume: $\rho(K)=|\mathcal{S}_{\text{reasoning}}^{(K)}|/|\mathcal{V}_{\text{semantic}}|.$
To substantiate our claim that this density grows exponentially, we propose the following assumption.
@@
(Exponential expressive redundancy of language.) For any non-trivial reasoning step $\xi_{l}$, the size of its semantic equivalence set $|\mathcal{V}(\xi_{l},k)|$, when realized by $k$ tokens, grows exponentially with $k$. That is, there exists a constant $c>1$ such that: $|\mathcal{V}(\xi_{l},k)|\geq c^{k}.$
This assumption is grounded in the combinatorial nature of language. Allowing even a few extra tokens creates an exponentially larger space of semantically equivalent expressions.
@@
(Continuum convergence of reasoning-level state space.) Under Assumption 2.9, in a $D$-dimensional semantic space, the reasoning-level state density, denoted as $\rho(K)$, increases exponentially with the reasoning token amount $K$, formally $\rho(K)=\Theta(c^{K})$. Let $s_{i}\bowtie s_{j}$ denotes that $s_{i},s_{j}$ are nearest neighbor of each other, the expected nearest neighbor distance of the semantic manifold is inversely related to the reasoning state density, formally $\mathbb{E}_{s_{i}\bowtie s_{j}}[\operatorname{dist}(s_{i},s_{j})]\propto\rho(K)^{-1/D}$.
The proof is provided in Appendix B.1. In Theorem 2.10, we established that the reasoning-level state space converges to a continuous semantic manifold in a macroscopic limit. Here, we further quantify the total “gap” between an ideal continuous update and the best possible discrete step, demonstrating that this gap vanishes as the total reasoning token amount $K$ increases.
[…]